Shift summary for Owl 11/12/99

Yee Bob Hsiung (hsiung@fnal.gov)
Fri, 12 Nov 1999 07:02:34 -0600

Shift Summary for Owl shift, 11/12/99

On Shift: Sasha Glazov, Bob Hsiung

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: MI, TeV, LINAC, Booster startup problems

========================================================================
Runs ended this shift: none

Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: Waiting for beam, running cosmic.
Pretty quiet most of the time except a flurry of
ktevana errors woke up shiftees and we could not
stop Elliott. Restart a new cosmic run.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

00:00 shift started with the same old message on Ch13:
"TeV B2 power supply problem, MI
ground fault. Est. 2-4 hours for beam."
continue cosmic run.
This could be a looooooooooooong night again.
00:45 Marry called from MCR to inform us that
they are still working on the problems and
no good estimate for beam.
02:58 MI ground fault and TeV B2 PS problems fixed,
but having peoblem with LINAC klystron vacuum,
est. 2 hrs for beam.
checking the cosmic run still running.
04:51 Ch13: Turning on in progress, est. 45 min.
Ooooooops! not so fast.
05:04 non-stop ktevana errors, such as illegal mod
num in stream 5, Ordinal number mismatch, etc.
can not be cured by repeated kfrends.
The readout is falling apart!
Stop and Restart a new cosmic run, the 8th INITs.
Ch13: Trouble with Booster correction element
PS. Experts en route.
Doze back to sleep!
06:00 DC gas check. Cool, 40 degreeF.
========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

When beam returns:

- Notify Bob T.
- Take muon run
- Flip NM4AN polarity to +
- Take pedestal run
- Update constants

========================================================================